The length of a square is increased by 8 and the width is decreased by 4. The area of the resulting is 12 square units more than the area of the square. How do I find the dimensions of the rectangle.
Answer by mananth(16946)

Let the length of side be x
area of square = x^2
side increased by 8 = x+8
side decreased by 4 = x-4
Resulting area = ( x-4)(x+8)
x^2+12= (x-4)(x+8)
x^2+12=x^2+8x-4x-32
x^2+12= x^2+4x-32
x^2-x^2-4x=-32-12
-4x=-44
x=11
x+8 is the length = 11+8=19
x-4= width = 11-4=7
The dimensions are 11 and 7